OpenStudy (word2):

Snowy's Snow Cones has a special bubble gum snow cone on sale. The cone is a regular snow cone that has a spherical piece of bubble gum nested at the bottom of the cone. The radius of the snow cone is 2 inches, and the height of the cone is 3 inches. If the diameter of the bubble gum is 0.5 inches, which of the following can be used to calculate the volume of the cone that can be filled with flavored ice? 1 over 3(3.14)(22)(3) − 4 over 3(3.14)(0.25^3) 1 over 3(3.14)(32)(2) − 4 over 3(3.14)(0.25^3) 1 over 3(3.14)(22)(3) − 4 over 3(3.14)(0.5^3) 1 over 3(3.14)(32)(2) − 4 over 3(3.14)(0.5^3)

OpenStudy (word2):

the answer is either 1 over 3(3.14)(22)(3) − 4 over 3(3.14)(0.253) or 1 over 3(3.14)(22)(3) − 4 over 3(3.14)(0.53) but i don't know which one

OpenStudy (phi):

the diameter of the bubble gum is 0.5 inches, the bubble gum makes a sphere. the volume of a sphere is \[ \frac{4}{3} \pi r^3\] where r is the radius of the sphere if the diameter is 0.5 inches, any idea what the radius is ?

OpenStudy (word2):

is it 1?

OpenStudy (word2):

@sammixboo

OpenStudy (word2):

@sweetburger

OpenStudy (sweetburger):

@word2 phi showed you that the diameter is .5 inches. The radius is always one half the diameter. Determine what r is\[\frac{ .5 }{ 2 }=r\]

OpenStudy (word2):

OOOOH okay thank you , i was super confused

OpenStudy (sweetburger):

yea np glad it clarified things.